# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@ 6.2 A Data Structure for $n$ Points
The overall query time is $O(n\frac{2^{p}}{\epsilon^{p+2}}\log\frac{n}{\delta}(\log\log n)^{p+3})$. Since we are running the preprocessing algorithm for all pairs of points, the space is $O(n^{2}\log d\frac{2^{p}}{\epsilon^{p+2}}\log\frac{n}{\delta}(\log\log n)^{p+3})$. ∎
7 $O(1)$-approximate sublinear representation of clusterings
Theorem 7.1.
Definition 7.2 (cost of clustering).
Consider a set $C$ of $k$ centers $c_{1},\dots,c_{k}$, and a set $P$ of points in $\mathbb{R}^{d}$ equipped with $\ell_{p}$-metric. The optimal cost of clustering is $\mathsf{optcost}(C,P)=\sum_{q\in P}\min_{c_{i}\in C}\|q-c_{i}\|_{p}^{p}$. The cost w.r.t. partition $\sigma:P\rightarrow[k]$ is $\mathsf{cost}(C,P,\sigma)=\sum_{q\in P}\|q-\sigma(q)\|_{p}^{p}$.
Corollary 7.3.
Consider a set $C$ of $k$ centers $c_{1},\dots,c_{k}\in[-\Delta,\Delta]^{d}$ equipped with $\ell_{1}$ or $\ell_{2}$ metric. We can efficiently construct a randomized data structure such that for any set $P\subset[-\Delta,\Delta]^{d}$ of points the following conditions hold. (i) The data structure reads $O(k\mathrm{polylog}(k,d,\Delta)/(\epsilon^{3}))$ coordinates of $P$. (ii) It returns a partition $\sigma:P\rightarrow[k]$ (iii) $\mathsf{cost}(C,P,\sigma)\leq O(1)\mathsf{optcost}(C,P)$. (iv) The data structure has a size of $O(k\mathrm{polylog}(k,d,\Delta)/(\epsilon^{3}))$ words.
3
4      for $t\in[T]$ do
5             for $b\in[d]$ do
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{n\choose 2}}\frac{\left|c_{i}^{(b)}-c_{j}^{(b)}\right|}{\|c_{i}-c_{j}\|_{1}}$;
7
8
10       Let $m\leftarrow O(\log(n)\log(\Delta d)/(\epsilon^{2}))$;
11       Sample $M\in\mathbb{R}^{[m]\times I}$, where each entry of $M$ follows the Cauchy distribution, whose density function is $c(x)=\frac{1}{\pi(1+x^{2})}$. Notice that $M(\cdot)\colon\mathbb{R}^{I}\rightarrow\mathbb{R}^{m}$ is an oblivious linear mapping [Ind06];
12
13      store $I$, $M(R)=\{Mr_{i}|i\in[n]\}$, $\{p^{(b)}|b\in I\}$, $M$
14Query $(q)$ // Inputs: $q\in[-\Delta,\Delta]^{d}$
15       Query $q^{(b)},\forall b\in I$;
16       Let $u\in\mathbb{R}^{I}$, where $\forall b\in I,u^{(b)}=q^{(b)}/p^{(b)}$;
18       Let $\hat{i}=\operatorname*{arg\,min}_{i\in[n]}F(Mr_{i},Mu)$;
19       return $\hat{i}$
Theorem 7.4.
$\mathbb{E}[\Gamma]=O(1)$.
Lemma 7.5.
It holds that $\mathbb{P}[G_{1}]\geq 1-\epsilon/(4\Delta d)$
Definition 7.6 (Good and bad regions).
Consider two points $c_{1},c_{2}\in\mathbb{R}^{d}$, and a query point $q\in\mathbb{R}^{d}$. We define the good region w.r.t. $(c_{1},c_{2},q)$ as
$\mathsf{Good}(c_{1},c_{2},q)=\{z\in[d]\mid\min\{c_{1}^{(z)},c_{2}^{(z)}\}-\frac{8}{\epsilon}|c_{1}^{(z)}-c_{2}^{(z)}|\leq q^{(z)}\leq\max\{c_{1}^{(z)},c_{2}^{(z)}\}+\frac{8}{\epsilon}|c_{1}^{(z)}-c_{2}^{(z)}|\}$
Lemma 7.7.
$\mathbb{P}[G_{2}]\geq 1-\epsilon/(4\Delta d)$
Proof.
Using Chernoff bound,
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}-\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right|\geq\epsilon_{2}\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right]\leq 2\exp\left(-\frac{\epsilon_{2}^{2}T\mathbb{E}[\sum_{z\in[d]}P_{t,z}]}{3(1+\frac{8}{\epsilon})\|c_{i}-c_{i^{*}}\|_{1}}\right)$
$=2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}{\|c_{i}-c_{i^{*}}\|_{1}}\right)\leq 2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon})}\cdot\frac{1}{2}\cdot\frac{13}{16}\right)\leq\frac{1}{2\Delta n},$
We conclude the lemma by union bound over all $c_{i}\neq c_{i^{*}}$ such that $(\|c_{i}-q\|_{1})/(\|c_{i^{*}}-q\|_{1})\geq 1+\epsilon$. ∎
Let $G=G_{1}\wedge G_{2}$. Lemmas 7.5 and 7.7, together with union bound gives the following.
Lemma 7.8.
$\mathbb{P}[G]\geq 1-\epsilon/(2\Delta d)$.
Lemma 7.9.
Proof.
Since we know that $\mathsf{dist}_{\mathsf{Bad}}(c_{i},q)-(1+\epsilon)\mathsf{dist}_{\mathsf{Bad}}(c_{i^{*}},q)\leq 0$, we have $\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}_{\mathsf{Good}}(c_{i^{*}},q)+\mathsf{dist}_{\mathsf{Bad}}(c_{i^{*}},q)}\geq 2\max\{1,\lambda\}$, which is $\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\|c_{i^{*}}-q\|_{1}}\geq 2\max\{1,\lambda\}$. Since $\mathsf{dist}_{\mathsf{Good}}(r_{i},u)\approx_{\epsilon}T\cdot\mathsf{dist}_{\mathsf{Good}}(c_{i},q)$, and $\|r_{i^{*}}-u\|_{1}=\lambda\cdot T\cdot\|c_{i^{*}}-q\|_{1}$, we have $\frac{\mathsf{dist}_{\mathsf{Good}}(r_{i},u)}{\|r_{i^{*}}-u\|_{1}}\geq\frac{2\max\{1,\lambda\}}{(1+\epsilon)\lambda}$. Thus $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 2/(1+\epsilon)\geq(1+\epsilon)^{2}$. ∎
We next bound the probability that the algorithm outputs a far center.
Lemma 7.10.
Proof.
By Lemma 7.9 it follows that $\mathsf{dist}(r_{i},u)>(1+\epsilon)^{2}\mathsf{dist}(r_{i^{*}},u)$. Since the event $G$ contains the event $G_{1}$ it follows that $F(Mr_{i},Mu)>F(Mr_{i^{*}},Mu)$. Thus, the algorithm does not output $i$. ∎
Lemma 7.11.
Proof.
$\leq 4\max\{\lambda,1\}$
The last inequality follows from Lemma 7.10. ∎
We use Lemma 7.11 to conclude the proof for the constant approximation factor.
Proof of 7.4.
$\leq\epsilon+4\mathbb{E}[\Lambda]+4\leq 8+\epsilon$ (11)
In Line 8 we use the fact that the maximal distance inside the $[-\Delta,\Delta]^{d}$ grid is $2\Delta d$ and Lemma 7.8. In Line 9 we use the fact $\mathbb{P}[\Lambda=\lambda\mid G]\leq\mathbb{P}[\Lambda=\lambda]/\mathbb{P}[G]$. In Line 10 we use Lemma 7.11. In Line 11 we use the fact that $\mathbb{E}[\Lambda]=1$. ∎
Claim 7.12.
With probability $1-1/(nd\Delta)$, we have $|I|\leq 2Tn$.
Proof.
Using Chernoff bound, we have
Thus, with probability $1/(nd\Delta)$, we have $|I|\leq 2E[|I|]\leq 2Tn$. ∎
Lemma 7.13.
Proof.
By 7.12, we have that the number of coordinates of $q$ that we query is $|I|=O(Tn)=O(\frac{n\log(n)\log(\Delta d)}{\epsilon^{3}})$. Next we analyze the space complexity of the data structure. Matrix $M$ is of wordsize $O(m|I|)$, $M(R)$ is of wordsize $O(mn)$, $I$ is of wordsize $O(|I|\log(d))$. Thus the space complexity of the our data strcuture is $O(m|I|+mn+|I|\log(d))=O(n\mathrm{polylog}(n,d,\Delta)/(\epsilon^{3}))$ wordsize. ∎
Appendix A Missing Proofs from Section 5
Claim A.1.
Proof.
From $g_{1}^{\prime}\approx_{\epsilon/16}g_{1}$ and $w^{\prime}\approx_{\epsilon/16}w$, we have $g_{1}^{\prime}-w^{\prime}\geq\left(1-\frac{\epsilon}{16}\right)g_{1}-\left(1+\frac{\epsilon}{16}\right)w.$ Let $\rho=\frac{3}{4}-\frac{\epsilon}{16}$. We rewrite $\frac{g_{1}}{w}\geq 1+\frac{\epsilon}{2}$ as $\left(1-\frac{\epsilon}{16}-\rho\right)g_{1}\geq\left(1+\frac{\epsilon}{16}-\rho\right)w$. This shows $\left(1-\frac{\epsilon}{16}\right)g_{1}-\left(1+\frac{\epsilon}{16}\right)w\geq\rho(g_{1}-w)$. Thus we have $g_{1}^{\prime}-w^{\prime}\geq\rho(g_{1}-w)$. ∎
Proof of 5.9.
Given $\frac{g_{1}+b_{1}}{g_{2}+b_{2}}\geq 1+\epsilon$, we start by deducting $1+\frac{\epsilon\delta^{\prime}}{8}$ on both sides: $\frac{g_{1}+b_{1}-(1+\frac{\epsilon\delta^{\prime}}{8})(g_{2}+b_{2})}{g_{2}+b_{2}}\geq 1+\epsilon-(1+\frac{\epsilon\delta^{\prime}}{8})\geq\frac{7}{8}\epsilon.$ Since $\frac{b_{1}}{b_{2}}\leq 1+\frac{\epsilon\delta^{\prime}}{8}$, we get $\frac{g_{1}-(1+\frac{\epsilon\delta^{\prime}}{8})g_{2}}{g_{2}+b_{2}}\geq\frac{7}{8}\epsilon.$ Using A.1 and letting $w=(1+\frac{\epsilon}{8})g_{2}$, we have $\frac{g_{1}^{\prime}-(1+\frac{\epsilon\delta^{\prime}}{8})g_{2}^{\prime}}{g_{2}+b_{2}}\geq\frac{21-\frac{7\epsilon}{4}}{32}\epsilon\geq\frac{5\epsilon}{8}.$ Since $g_{2}^{\prime}+b_{2}^{\prime}\leq\frac{1}{\delta^{\prime}}(g_{2}+b_{2})$, we get $\frac{g_{1}^{\prime}-(1+\frac{\epsilon\delta^{\prime}}{8})g_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{5\epsilon\delta^{\prime}}{8}.$ Rewriting the left-hand side as $\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}-\frac{\epsilon\delta^{\prime}}{8}(g_{2}^{\prime}+b_{2}^{\prime})-b_{1}^{\prime}+(1+\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{5\epsilon\delta^{\prime}}{8}.$ Rearranging terms gives $\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}-b_{1}^{\prime}+(1+\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{5\epsilon\delta^{\prime}}{8}+\frac{\epsilon\delta^{\prime}}{8}.$ Using $\frac{b_{1}^{\prime}}{b_{2}^{\prime}}\geq 1-\frac{\epsilon\delta^{\prime}}{8}$, we get $\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}+(\frac{\epsilon\delta^{\prime}}{4})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}=\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}-(1-\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}+(1+\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{3\epsilon\delta^{\prime}}{4}.$
Appendix B Implication for Range Search
@@
Definition B.1 ($(1+\epsilon)$-approximate range search problem).
@@ 2       Let $I\leftarrow\emptyset$ be a multiset and $T\leftarrow O(\log(n/\delta)/\epsilon)$;
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{n\choose 2}}\frac{\left|c_{i}^{(b)}-c_{j}^{(b)}\right|}{\|c_{i}-c_{j}\|_{1}}$;
@@ 16                   Add $i$ to $W$;
This is a contradiction to the choice of $c_{i}$. ∎
@@
$\left(1-\frac{1}{e}\right)\frac{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})}{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})+\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}$ $\geq\left(1-\frac{1}{e}\right)\left(\frac{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})+\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})}\right)^{-1}$
@@
Since we repeat for $T=O(\log(n/\delta)/\epsilon)$ iterations, we sample a coordinate from $\mathsf{Bad}$ with probability at least $1-\delta/n$. Using union bound we succeed for all points $i\in[n]$ with probability at least $1-\delta$
Appendix C Reducing Approximate Nearest Neighbor under $\ell_{2}$-norm to Feature Selection for $k$-means Clustering
@@ 18       Query $t$ coordinates of $q$ to compute $y^{\intercal}=q^{\intercal}\Omega S$;
C.1 Preliminaries
Before proving the Theorem C.1, we introduce some notations.
@@ C.2 Main Proof
To calculate the eigenvalues of $B^{\intercal}B$, we let $\lambda I-B^{\intercal}B=0$. After some manipulation[Gol07], we know the eigenvalues $\lambda$ of $B^{\intercal}B$ corresponds to the solutions of
@@
Thus for any $10<\epsilon<$ 1, there exists sufficiently large $w$, such that $\|P-PV_{n}(C)V_{n}(C)^{\intercal}\|_{F}^{2}\leq(1+\epsilon)h(w)$, i.e. $\|B-BV_{n}(C)V_{n}(C)^{\intercal}\|_{F}^{2}\leq(1+\epsilon)\|B-B_{n}\|_{F}^{2}$.
Definition C.3.
@@
Theorem C.4.
(Theorem 11 and Lemma 4 of [BZMD15]) Let $B\in\mathbb{R}^{m\times d}$ be a set of $m$ points in $\mathbb{R}^{d}$, and $k$ be the input of k-means clustering. Let $Z\in\mathbb{R}^{d\times k}$ be an approximation of $V_{k}(B)$, such that $Z^{\intercal}Z=I_{k\times k},(B-BZZ^{\intercal})Z=\mathbf{0}_{m\times k}$, and $\|B-BZZ^{\intercal}\|_{F}^{2}\leq(1+\epsilon)\|B-B_{k}\|_{F}^{2}$. Then let $[\Omega,S]=SelectFeature(Z,t)$, with $t=O(k\log k/\epsilon^{2})$. Let $B^{\prime}=B\Omega S\in\mathbb{R}^{m\times t}$ be the points after feature selection. Then a $\gamma$-approxmation partition $\rho$ for $B^{\prime}$ is $(1+(2+\epsilon)\gamma)$-approximation parition for $B$ with probability $0.2$.
Theorem C.5.
(Theorem 1.3 of [MMR19]) Let $B^{\prime}\in\mathbb{R}^{m\times t}$ be a set of $m$ points in $\mathbb{R}^{t}$ and $k$ be the input of $k$-means clustering. Given $40<\epsilon,\delta<1/$ 4. Let $r=\log(k/\epsilon\delta)/\epsilon^{2}$. There exists a family of random maps $\pi:\mathbb{R}^{t}\rightarrow\mathbb{R}^{r}$, which is independent of $B^{\prime}$, such that $\underset{M\sim\pi}{\mathbb{P}}\left[\forall\rho,cost(B^{\prime},\rho)\approx_{\epsilon}cost(M(B^{\prime}),\rho)\right]\geq 1-\delta$.
Proof.
(for Theorem C.1) Recall that $B$ can be viewed as a point set of size $(wn+1)$ in $\mathbb{R}^{d}$. Let $B^{\prime}=B\Omega S$ as defined in Theorem C.4, which can be viewed as a point set of size $(wn+1)$ in $\mathbb{R}^{t}$. Let $M$ be the random dimension reduction map in Theorem C.5. Then $M(B^{\prime})$ can be viewed as a point set of size $(wn+1)$ in $\mathbb{R}^{r}$.
@@
From Theorem C.5, we know that with probability $1-\delta$,
@@
Combine them gives $\textsf{cost}(B^{\prime},\rho_{2})\leq(1+\epsilon)\textsf{cost}(M(B^{\prime}),\rho_{2})\leq(1+\epsilon)(1+\epsilon)\textsf{cost}(M(B^{\prime}),\rho_{1})\leq(1+\epsilon)^{2}\textsf{cost}(B^{\prime},\rho_{2})$. This shows that $\rho_{2}$ is $(1+\epsilon)^{2}$-approximation for $B^{\prime}$ with probability $1-\delta$. Furthermore, Theorem C.4 tells us that $\rho_{2}$ is $(3+10\epsilon)$-approximation for $B$ with probability $0.2-\delta$, i.e.
@@
The reason that we can not use the same trick in [CEM+15] to achieve $(1+\epsilon)$-ANN is, it select features based on $Z=V_{n}(C)$ and $B-BZZ^{\intercal}=\begin{bmatrix}0\\ 0\\ \vdots\\ 0\\ P-PZZ^{T}\end{bmatrix}$, which requires some information of the query.
Appendix D Lower Bounds
D.1 Lower Bound for Approximate Nearest Neighbor

### v2

@@ D.2 A Data Structure for $n$ Points
The overall query time is $O(n\frac{2^{p}}{\epsilon^{p+2}}\log\frac{n}{\delta}(\log\log n)^{p+3})$. Since we are running the preprocessing algorithm for all pairs of points, the space is $O(n^{2}\log d\frac{2^{p}}{\epsilon^{p+2}}\log\frac{n}{\delta}(\log\log n)^{p+3})$.
Appendix E Implication for Range Search
@@
Definition E.1 ($(1+\epsilon)$-approximate range search problem).
@@ 2       Let $I\leftarrow\emptyset$ be a multiset and $T\leftarrow O(\log(n/\delta)/\epsilon)$;
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{\binom{n}{2}}}\frac{\left|c_{i}^{(b)}-c_{j}^{(b)}\right|}{\|c_{i}-c_{j}\|_{1}}$;
@@ 16                   Add $i$ to $W$;
This is a contradiction to the choice of $c_{i}$.
@@
$\left(1-\frac{1}{e}\right)\frac{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})}{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})+\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}$
$\geq$ $\left(1-\frac{1}{e}\right)\left(\frac{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})+\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}{\mathsf{dist}_{\mathsf{Bad}}(c_{i},c_{i^{*}})}\right)^{-1}$
@@
Since we repeat for $T=O(\log(n/\delta)/\epsilon)$ iterations, we sample a coordinate from $\mathsf{Bad}$ with probability at least $1-\delta/n$. Using union bound we succeed for all points $i\in[n]$ with probability at least $1-\delta$.
Appendix F Reducing Approximate Nearest Neighbor under $\ell_{2}$-norm to Feature Selection for $k$-means Clustering
@@ 18       Query $t$ coordinates of $q$ to compute $y^{\intercal}=q^{\intercal}\Omega S$;
F.1 Preliminaries
Before proving the Theorem F.1, we introduce some notations.
@@ F.2 Main Proof
To calculate the eigenvalues of $B^{\intercal}B$, we let $\lambda I-B^{\intercal}B=0$. After some manipulation[21], we know the eigenvalues $\lambda$ of $B^{\intercal}B$ corresponds to the solutions of
@@
Thus for any $10<\epsilon<$ 1, there exists sufficiently large $w$, such that $\|P-PV_{n}(C)V_{n}(C)^{\intercal}\|_{F}^{2}\allowbreak\leq(1+\epsilon)h(w)$, i.e. $\|B-BV_{n}(C)V_{n}(C)^{\intercal}\|_{F}^{2}\leq(1+\epsilon)\|B-B_{n}\|_{F}^{2}$.
Definition F.3.
@@
Theorem F.4.
(Theorem 11 and Lemma 4 of [11]) Let $B\in\mathbb{R}^{m\times d}$ be a set of $m$ points in $\mathbb{R}^{d}$, and $k$ be the input of k-means clustering. Let $Z\in\mathbb{R}^{d\times k}$ be an approximation of $V_{k}(B)$, such that $Z^{\intercal}Z=I_{k\times k},(B-BZZ^{\intercal})Z=\mathbf{0}_{m\times k}$, and $\|B-BZZ^{\intercal}\|_{F}^{2}\leq(1+\epsilon)\|B-B_{k}\|_{F}^{2}$. Then let $[\Omega,S]=SelectFeature(Z,t)$, with $t=O(k\log k/\epsilon^{2})$. Let $B^{\prime}=B\Omega S\in\mathbb{R}^{m\times t}$ be the points after feature selection. Then a $\gamma$-approxmation partition $\rho$ for $B^{\prime}$ is $(1+(2+\epsilon)\gamma)$-approximation parition for $B$ with probability $0.2$.
Theorem F.5.
(Theorem 1.3 of [29]) Let $B^{\prime}\in\mathbb{R}^{m\times t}$ be a set of $m$ points in $\mathbb{R}^{t}$ and $k$ be the input of $k$-means clustering. Given $40<\epsilon,\delta<1/$ 4. Let $r=\log(k/\epsilon\delta)/\epsilon^{2}$. There exists a family of random maps $\pi:\mathbb{R}^{t}\rightarrow\mathbb{R}^{r}$, which is independent of $B^{\prime}$, such that .
Proof F.6.
(for Theorem F.1) Recall that $B$ can be viewed as a point set of size $(wn+1)$ in $\mathbb{R}^{d}$. Let $B^{\prime}=B\Omega S$ as defined in Theorem F.4, which can be viewed as a point set of size $(wn+1)$ in $\mathbb{R}^{t}$. Let $M$ be the random dimension reduction map in Theorem F.5. Then $M(B^{\prime})$ can be viewed as a point set of size $(wn+1)$ in $\mathbb{R}^{r}$.
@@
From Theorem F.5, we know that with probability $1-\delta$,
@@
Combine them gives $\textsf{cost}(B^{\prime},\rho_{2})\leq(1+\epsilon)\textsf{cost}(M(B^{\prime}),\rho_{2})\leq(1+\epsilon)(1+\epsilon)\textsf{cost}(M(B^{\prime}),\rho_{1})\leq(1+\epsilon)^{2}\textsf{cost}(B^{\prime},\rho_{2})$. This shows that $\rho_{2}$ is $(1+\epsilon)^{2}$-approximation for $B^{\prime}$ with probability $1-\delta$. Furthermore, Theorem F.4 tells us that $\rho_{2}$ is $(3+10\epsilon)$-approximation for $B$ with probability $0.2-\delta$, i.e.
@@
The reason that we can not use the same trick in [17] to achieve $(1+\epsilon)$-ANN is, it select features based on $Z=V_{n}(C)$ and $B-BZZ^{\intercal}=\begin{bmatrix}0\\ 0\\ \vdots\\ 0\\ P-PZZ^{T}\end{bmatrix}$, which requires some information of the query $q$ in the preprocessing of the nearest neighbor data structure.
Appendix G Lower Bounds
G.1 Lower Bound for Approximate Nearest Neighbor